OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

OpenStudy (anonymous):

x/6 +y/5 =1 slope intercept form is y=-5/6x +5

OpenStudy (anonymous):

points are (6,0),(0,5) m=5/-6 y=mx+c y=-5/6x+5